OpenStudy (anonymous):

Rewrite the expression with positive exponents: 1/(3x-3)

OpenStudy (anonymous):

it already has a positive exponent

OpenStudy (anonymous):

i knooow thats what i dont get

OpenStudy (phi):

is that supposed to be \[ \frac{1}{3x^{-3} } \]

OpenStudy (phi):

or \[ \frac{1}{(3x)^{-3}} \]

OpenStudy (phi):

or \[ \frac{1}{3x-3} \]

OpenStudy (anonymous):

i cant ight it like that

OpenStudy (phi):

which one is it? post 1, post 2 or post 3?

OpenStudy (anonymous):

neither the answer has to be right in a box a thin box

OpenStudy (phi):

The posts are meant to be the question, not the answer. I am not sure what your question is

OpenStudy (anonymous):



OpenStudy (anonymous):

is that what you mean?

OpenStudy (anonymous):